A random sample of 46 college students reported the number of hours per day they typically spend on social media. Their sample mean is, M = 2.61, and their standard deviation is, s = 1.02. If you plan to construct a 95% confidence interval for the population mean, what t critical value should you use based on this data?
Answer #1

Here,

Sample size(n)=46

Hence,

Degrees of freedom=n-1=46-1=45

And required critical t value is

critical value=2.0141

Thank you.